OpenStudy (anonymous):

A circle has a diameter of twenty yards. What is its area? Use three point one four for pi one hundred twenty-five point six square yards three hundred fourteen square yards two hundred fifty-one point two square yards six hundred twenty-eight square yards

OpenStudy (anonymous):

diameter of circle d= 20 yards radius r = 20/2 = 10 yards area of circle \[\huge \pi r^2 = 3.14 \times 10^2 = 3.14 \times 100 = 314 \] Hence area of the circle is 314 yards

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Hence option three hundred fourteen square yards is the correct answer...